OpenStudy (anonymous):

4.An ice cream store is making waffle cones. The finished cone should have a radius of 1 inch and be 5 inches tall. The slant height of the cone is 5.1 inches. How much waffle will they need to make each cone? Express your answer to the nearest square inch. A. 6 in2 B. 16 in2 C. 18 in2 D. 19 in2

Directrix (directrix):

SA = πr2 + πrl Where, r is the radius l is the slant height We don't need the area of the base. It is open. LA = pi * r * l where LA is lateral area LA = pi * 1 * 5.1 LA = 16.02 B. 16 in2 --> ANS